OpenStudy (anonymous):

What is the 10th term of the sequence 81, 27, 9,...

OpenStudy (anonymous):

A. 1/179 B. 1/243 C. 1/81 D. 1/810

OpenStudy (kc_kennylau):

What's your answer?

OpenStudy (anonymous):

I think it's D

OpenStudy (kc_kennylau):

Nope it isn't D

OpenStudy (anonymous):

Hmm...let me work it again I might have gone to far

OpenStudy (anonymous):

C. 1/81?

OpenStudy (kc_kennylau):

Nope

OpenStudy (anonymous):

Wait that's 9th term it's 1/243

OpenStudy (mathmale):

I strongly recommend that you find the appropriate expression for the nth term of this geometric sequence.

OpenStudy (kc_kennylau):

Yep that's correct

OpenStudy (mathmale):

What is the initial value, a? What is the common ratio, r?

OpenStudy (kc_kennylau):

By the way do you understand the concept?

OpenStudy (anonymous):

Yea it was division by 3

OpenStudy (mathmale):

Or, "our common ratio is 1/3."

OpenStudy (mathmale):

so, r = 1/3. What is the first term, a? a = ?

OpenStudy (anonymous):

a = 81 (from the sequence it was given)

OpenStudy (mathmale):

Right. Could you now develop a formula for the nth term of this geometric sequence? \[a _{n+1}= ?\]

OpenStudy (anonymous):

a(1/3) + 1 = ? (would I just put 81 then 27 then 9 etc.. in the a term?)

OpenStudy (kc_kennylau):

How would you find the next term of this sequence? Describe the process by words please.

OpenStudy (mathmale):

What we need here is a formula in n (that means that n is your independent variable) that correctly predicts the nth term of this geometric sequence. Your tentative result does not involve n, so is not correct. Look: If n=0, a*(1/3)^0=81*(1/3)^0 = ?

OpenStudy (anonymous):

81*(1/3)^0 would equal 81?

OpenStudy (anonymous):

ok ... 81/3 = 27 27/3 = 9 Do that a few more times.................. its going to get negative.

OpenStudy (kc_kennylau):

There are two presentations for a formula of a geometric sequence

OpenStudy (anonymous):

or... Arithmetic sequence:1, 3, 5, 7 Notice how I add 2 everytime. This is arithmetic.1, 2, 4, 8 This is geometric. I am not adding or subtracting the same number every time, as 1 to 2 is +1, while 2 to 4 is +2. So what is the pattern? It is geometric. I am multiplying by 2 to get my next sequence. So in your problem, you are not adding or subtracting the same number every time because clearly this is not the case. So this is NOT an arithmetic sequence, but if you can find a common multiplying or dividing factor you can identify it as geometric. As others have mentioned, you are dividing by 3 (or multiplying by 1/3) each time to obtain your next time.

OpenStudy (kc_kennylau):

One is called recursive formula, which is recursive (thank you captain obvious)

OpenStudy (kc_kennylau):

One is called explicit formula or closed-form expression, which is not recursive.

OpenStudy (anonymous):

So this would be an explicit formula?

OpenStudy (kc_kennylau):

What do you mean by this?

OpenStudy (anonymous):

Nvm XD Thought you where wanting me to choose which formula to use

OpenStudy (kc_kennylau):

An explicit formula for a geometric sequence would be A_n = a * r^(n-1)

OpenStudy (kc_kennylau):

Or in LaTeX, \(\Large A_n=a\cdot r^{n-1}\).

OpenStudy (kc_kennylau):

Where \(A_n\) is the n-th term, a is the first term, and r is the common ratio

OpenStudy (mathmale):

81*(1/3)^0 would equal 81? YES. What about 81*(1/3)^1? What about 81*(1/3)^2? You'll see, through experimentation, that this fits Kenny's explicit formula perfectly.

OpenStudy (kc_kennylau):

A general recursive formula for a geometric sequence would be A_(n+1) = r * A_n

OpenStudy (kc_kennylau):

Or in LaTeX, \(\Large A_{n+1}=r\cdot A_n\).

OpenStudy (mathmale):

If you agree, take Kenny's formula and substitute n=10. What's the 10th term?

OpenStudy (kc_kennylau):

Where \(A_{n+1}\) is the next term, \(A_n\) is the previous term, and \(r\) is the common ratio.

OpenStudy (mathmale):

Also, @Cpt.Sparz, would you look at Kenny's two different formulas and decide which one would be better suited for use in predicting the 10th term of this geom. seq.?

OpenStudy (mathmale):

And why?

OpenStudy (anonymous):

(explicit formula) 81*1/3^10 - 1 = 1/243

OpenStudy (anonymous):

Explicit formula comes easier to me

OpenStudy (mathmale):

That'd be my choice also. Note that your response would be better if written as\[81*(1/3)^{10-1} = 1/243\]

OpenStudy (anonymous):

Oh ok...I don't know how to work that :3

OpenStudy (anonymous):

Like how to make it look that...

OpenStudy (mathmale):

I'd suggest learning how to use Equation Editor, if you haven't already. I'd choose the explicit formula because we need only substitute n=10 there, whereas in the case of the recursion formula, you have to calculate all 10 terms of the sequence.

OpenStudy (kc_kennylau):

You only have to add in parentheses.

OpenStudy (kc_kennylau):

81*1/3^(10 - 1) = 1/243 would be so much better than 81*1/3^10 - 1 = 1/243

OpenStudy (kc_kennylau):

Don't have to use the equation editor so often :p

OpenStudy (anonymous):

Oh ok thanks @kc_kennylau @mathmale

OpenStudy (anonymous):

\[81 * (\frac{ 1 }{ 3 })^{10 - 1} = \frac{ 1 }{ 243 }\]

OpenStudy (kc_kennylau):

That's better

OpenStudy (mathmale):

Note that \[81*(1/3)^{10-1} = 1/243\] shows that you're on the right track, but 1/243 is not the correct 10th term. \[81*(1/3)^{10-1} = 81*(\frac{ 1 }{ 3 })^{9}=\frac{ 3^4 }{ 3^9 }=??\]

OpenStudy (mathmale):

Note that 81=3^4.

OpenStudy (mathmale):

How would you reduce \[\frac{ 3^4 }{ 3^9 }?\]

OpenStudy (anonymous):

\[\frac{ 3^{4} }{ 3^{9} } = \frac{ 1 }{ 243 }\]

OpenStudy (mathmale):

Right. Sorry, I was wrong; 1/243 was correct all along. But this gives you an alternative way of evaluating the 10th term.

OpenStudy (anonymous):

Yea! I've never seen this way before, but I like it...more neat!

OpenStudy (mathmale):

cool. Hope to work with you again sometime soon! Bye for now.

OpenStudy (anonymous):

Thanks again!